OpenStudy (anonymous):

7. When Katie ordered her wedding dress she had to pay a deposit equal to two-thirds of the full price of the dress. She paid a $500 deposit. What is the full price of her wedding dress? Write the equation then solve. A. B. C. D.

hartnn (hartnn):

let full price be x $ so (2/3) x = 500 x= ?

OpenStudy (anonymous):

A.2/3d=500;d=333.33 B.d/3=500;d=1,500 C.3/2d=500;d=333.33 D.2/3d=500;d=750

hartnn (hartnn):

(2/3) d = 500 d= ?

OpenStudy (anonymous):

750?

OpenStudy (anonymous):

or 333.33?

OpenStudy (anonymous):

im so lost

hartnn (hartnn):

750 is correct

OpenStudy (anonymous):

so its D

hartnn (hartnn):

yup